Marigold Company purchased equipment for $1017000 on January 1, 2017, and will use the double-declining-balance method of depreciation. It is estimated that the equipment will have a 3-year life and a $36000 salvage value at the end of its useful life. The amount of depreciation expense recognized in the year 2019 will be

a. $77000.

b. $85880.

c. $121880.

d. $113000.
Answer #1

The correct answer A) $77000

Caluclations

Firstly we have to calculate the rate of double declining depreciation rate

= 1/ life *2*100

= 1/3*2*100

= 66.67%

WDVas on 1 January = cost * (1- depreciation rate)* (1- depreciation rate)

= 1017000*(1-66.67%)*(1-66.67%)

=$ 113000

Since, the year 2019 is the last useful year, so asset has to be depreciated to its salvage value in this year

Depreciation for year 2019 = WDV as on 1 january 2019- salvage value

= 113000-36000

= $ 77000

Thus the correct answer is $ 77000